OpenStudy (anonymous):

The following problem: \[h(x)=7x^2 - 5x -2\] Factors into this: \[7(x-1)(x+2/7)\] The problem can be solved using the quadratic formula: \[ax^2 +bx +c = 0\] \[x=(-b+-\sqrt{b^2-4ac)/})2a\] However, can the problem also be solved using the following formula? \[x2+px+q=0\] \[x=-p/2 +- \sqrt{(p/2)^2 - q}\]

mathslover (mathslover):

\[\large{h(x) = 7x^2-5x-2}\] \[\large{h(x) = 7x^2+7x-2x-2}\] \[\large{h(x) = 7x(x+1)-2(x+1)}\] \[\large{h(x) = (7x-2)(2x+1)}\]

OpenStudy (jiteshmeghwal9):

\[7x^2-7x+2x-2\]\[7x(x-1)+2(x-1)\]\[(7x+2)(x-1)\]

OpenStudy (anonymous):

I can solve the problem, but what I need is to udnerstand the formula x2+px+q=0. So far I haven't been able to solve the problem using this particular formula. :/

OpenStudy (jiteshmeghwal9):

take common\[7(x-1)(x+2/7)\]

OpenStudy (anonymous):

Still not using the formula though. I need to solve it using: \[x2+px+q=0\] \[x=-p/2 +- \sqrt{(p/2)^2 - q}\]

mathslover (mathslover):

@math0101 it is NOT given to us that : h(x) = 0 hence we can't use that formula